HETAV PATEL

Grade 11,

X²-X(1-a)-(a+2)=0, find the integral values of a for which the quadratic equation has integer roots

X²-X(1-a)-(a+2)=0, find the integral values of a for which the quadratic equation has integer roots

Grade:11

1 Answers

Meet
6 years ago
For integer roots Discriminant must be a perfect square and also check out the values by putting in Shree dharacharya method as D = a^2+2a+9= (a+1)^2 +8 now by this a=0,a=-2.Now if a=0 then roots are (2,-1)And if a=-2 then roots are (3,0).